# Autoformalization in the Era of Large Language Models: A Survey

### v1

@@ Abstract
DeepSeek-Prover-V1 miniF2F1,2 Autoformalized proof, Data construction
LeanEuclid LeanEuclid2,*, Elements 1, UniGeo1 Autoformalized proof, Evalute Autoformalization
Isa-AutoFormal Math1, miniF2F1 Evalute Autoformalization, Informalization
Retrieval Augmented autoformalization IsarMathLib2 Retrieval Augmented
DeepSeekMath GSM8K1, MATH1, SAT1, OCW Courses1, MGSM-zh1, CMATH1, Gaokao-MathCloze1, Gaokao-MathQA1, miniF2F1 Autoformalized proof
DRAFT, SKETCH, AND PROVE miniF2F1 DSP
FORMALALIGN miniF2F1, FormL41,2, MMA2 Evalute Autoformalization
GFLEAN textbook Mathematical Proofs1 AST simplification and transformation
Lean Workbook miniF2F2, ProofNet2 Data construction
LEAN-GitHub miniF2F1, ProofNet1, Putnam1 Data construction
LeanDojo LeanDojo1,2,* Retrieval Augmented, Autoformalized proof
LEGO-PROVER miniF2F1,2,* Retrieval Augmented, Autoformalized proof
LLEMMA Math1, GSM8K1, miniF2F1, ProofFile2,* Autoformalized proof
LYRA miniF2F1,2 Autoformalized proof, DSP, Evalute Autoformalization
MMA miniF2F1, ProofNet1, MMA2,* Informalization, Data construction
Pantograph miniF2F1 Autoformalized proof, Data construction, DSP
PDA FormL41,2,* Data construction, Informalization
ProofNet ProofNet1,2,* Retrieval Augmented, Informalization
SubgoalXL miniF2F1,2 Autoformalized proof
Herald miniF2F1,Extract Theorem1,College CoT1 Retrieval Augmented,Data construction,Autoformalization
DeepSeek-Prover-V2 miniF2F1,ProofNet1, Putnam1,CombiBench1,ProverBench1,* Autoformalized proof,Retrieval Augmented,Data construction,DSP
Kimina-Prover miniF2F1 Autoformalized proof,Autoformalization,Data construction
Goedel-Prover miniF2F1, Putnam1,ProofNet1,Lean Workbook1,2 Autoformalized proof
STP miniF2F1, Putnam1,ProofNet1 Autoformalized proof
Leanabell-Prover miniF2F1 Autoformalized proof,Data construction
BFS-Prover miniF2F1 Autoformalized proof,Retrieval Augmented,Data construction
InternLM2.5-StepProver miniF2F1,Putnam1,ProofNet1 ,LeanWorkbook-Plus1,* Autoformalized proof,Data construction
Table 2: Datasets with code repos link and description
Dataset Size Task Description
miniF2F 488 Theorem Proving statements from olympiads as well as high-school and undergraduate maths classes
LeanEuclid 173 Autoformalization Euclidean geometry problems manually formalized in Lean
Math 12,500 Fine-tuning 12,500 challenging competition mathematics problems
FORML4 17,137 Pre-training,Autoformalization informalize theorems extracted from Mathlib 4
ProofNet 374 Autoformalization, Theorem Proving The problems are drawn from popular undergraduate pure mathematics textbooks
PutnamBench 657 Theorem Proving competition mathematics problems sourced from the William Lowell Putnam Mathematical Competition
LeanDojo 122,517 Theorem Proving theorems,proofs,tactics,premises from mathlib4
GSM8K 8,792 Question Answering high quality linguistically diverse grade school math word problems
miniCTX 762 Theorem Proving theorems sourced from real Lean projects and textbooks
arxiv2Formal 50 Autoformalization theorems from arXiv papers formalized in Lean
DeepSeek-Prover-V1 27,503 Theorem Proving Lean4 proof data derived from high-school and undergraduate-level mathematical competition problems
Lean-Workbook 13,517 Autoformalization,Theorem Proving collection of mathematical problems and their proof information
Lean-GitHub 20,446 Autoformalization,Theorem Proving theorems with formal proofs and tactics from 2,133 files
Compfiles 205 Theorem Proving A collection of olympiad-style math problems and their solutions, formalized in Lean 4
Herald 624,436 Autoformalization,Theorem Proving The dataset generated from Herald pipeline on Mathlib4, containing 580k valid statements and 44k NL-FL proofs
ProverBench 325 Theorem Proving A benchmark dataset containing 325 formalized problems to advance neural theorem proving research, including 15 from the prestigious AIME competitions (years 24-25)
CombiBench 100 Theorem Proving A manually produced benchmark, including 100 combinatorial mathematics problems of varying difficulty and knowledge levels.
STP_Lean_0320 3,262,558 Theorem Proving A dataset including mathlib4 examples, LeanWorkbook proofs, and self-play conjecture proofs
Leanabell-Prover-Formal-Statement 1,133,283 Theorem Proving A dataset including released data by existing models and synthetic data from informal math problems.
@@ 1 Introduction
Beyond mathematics, however, the core ideas of autoformalization—translating informal natural language reasoning into precise, verifiable representations—hold broader significance. In domains such as scientific hypothesis generation, legal analysis, or even multi-hop question answering, ensuring the internal consistency and correctness of model-generated reasoning is increasingly important. Autoformalization provides a promising mechanism for grounding LLM reasoning in formal logic, enabling systematic checking of inference steps and enhancing the verifiability of LLM outputs. As LLMs move toward solving problems that demand not just fluent answers but reliable reasoning, autoformalization will play a crucial role in closing the gap between linguistic generation and logical soundness.
@@
Beyond mathematics, however, the core principles of autoformalization—translating informal or semi-formal language into precise, verifiable representations—hold broader significance. As LLMs are increasingly applied to tasks like algorithm design, code generation, and system specification Shi et al. (2023); Di et al. (2024), ensuring that their outputs are not only syntactically correct but also semantically sound becomes critical. Autoformalization offers a promising mechanism for bridging this gap: by converting natural language specifications, pseudo-code, or even informal correctness claims into formal logic or verifiable code contracts, it enables rigorous checking of algorithmic behavior. In domains such as software engineering, security auditing, and scientific computing, this capability could substantially enhance the trustworthiness and reliability of LLM-generated artifacts. Viewed from this perspective, autoformalization is not merely a tool for mathematical proof—it is a foundational enabler for building LLM systems that can reason, code, and verify in an integrated and principled way.
@@
This survey provides a comprehensive and structured review of the recent advances in automatic formalization, focusing on the current applications in mathematical theorem proving and the possibility of combining automatic formalization with formal verification of large models to generate fully trustworthy outputs in the future. §2 shows a review of the historical and technological foundations of autoformalization. In §3, we dissect the autoformalization workflow, covering key components such as data preprocessing, modeling strategies, postprocessing techniques, and evaluation protocols, and also dissect state of the art model approaches to mathematical theorem proving. In §4, we conduct a deep analysis of the involved mathematical problems from the mathematical problem domain, the mathematical problem difficulty level and the mathematical problem abstraction level. In §5, we explore the connection between automatic formalization and the formal verification of large model output from multiple fields. We summarize notable open-source models and datasets that support current research in §6; In §7, we discuss major open challenges and promising future research directions. Finally, §8 compares our survey with existing related surveys, clarifying its unique perspective and contributions.
@@ 2.2 Autoformalization in LLM Output Formal Verification
To address this challenge, recent research has explored the use of autoformalization as a means of translating natural language outputs from LLMs into precise, verifiable representations using formal systems such as first-order logic Ryu et al. (2024) or arithmetic system such as Peano arithmeticPeano (1889). In this context, autoformalization acts as a bridge between the inherent ambiguity of natural language and the strict requirements of formal verification. By enabling symbolic reasoning over LLM outputs, autoformalization enhances correctness, improves reliability, and promotes the safe and trustworthy deployment of LLMs in tasks requiring rigorous reasoning and high assurance.
@@ 3.1.2 Increase data quantity
Due to a large number of written by human still didn’t get the full use of formal language corpus, in order to solve this problem Wu et al. (2024b) collects almost all of the Lean4 repositories on GitHub, and makes suitable training data, increases the amount of data, and experiments show that training on such a large scale dataset, The model improves performance in a variety of mathematical problem domains.
@@ 3.2.2 LLM-based Autoformalization
In natural language machine translation related work, the quality of translation largely depends on the quality of parallel data between the two languages. But such parallel data are rare and difficult to manage, and back-translation is one of the most effective ways to improve the quality of translation. Azerbayev et al. (2023) use OpenAI distillation of the Codex model of back translation, in order to improve the model of automatic formal ability. Jiang et al. (2024) They used GPT-4 to generate informal language data from an existing library of formal languages. They have empirically and analytically concluded that informal is much simpler than formal, so back-translation is part of the Autoformalization effort.
@@ 3.2.4 Evaluatation
In the Autoformalization workflow, syntactic accuracy and semantic equivalence at the evaluation level are key factors to ensure the success of the autoformalization process. Syntactic accuracy refers to the ability to correctly translate into formal language, that is, whether the final formal result can pass the type check of Lean, Coq and other environments. The type checking of Lean, Coq and other environments does not mean that it is completely correct, because the formalized results also need to be semantically equivalent to the corresponding natural language version, so it is equally important to evaluate the semantic equivalence. The purpose of grammatical accuracy is to ensure the result of autoformalization through the formal language proof assistant’s type checking mechanism, which mainly involves two aspects: 1) at the level of code generation: the generated formal language code has no grammatical errors; 2) At the level of mathematical reasoning: the generated formal language code has a correct logical chain. The most common way to assess accuracy is to compare automatically generated formal expressions with the results of expert manual formalization. Such evaluation often rely on recall metrics or manual inspections(Wu et al. (2022) Agrawal et al. (2022), Azerbayev et al. (2023), Jiang et al. (2023a))to judge the consistency of the two in mathematical content. Although it is right, this method is time-consuming and subjective, so it is commonly used to calibrate automatic evaluate tools. For the accuracy of mathematical reasoning, such as whether the model corresponds to the specific meaning of each variable, or whether the proof strategy used in the proof process is correct, For the former, Li et al. (2024a)suggests the method of symbolic equivalence . Split statements into premises and conclusions (assuming that premises do not contradict themselves) When the premises and conclusions of the formal results to be verified are respectively equivalent to the corresponding correct premises and conclusions by ATP proof, it is explained that the model corresponds to the specific meaning of each variable. For the latter Zheng et al. (2024) leverage the feedback to evaluate whether the proof strategy is correct and the tool correction mechanism is used to resolve it. For semantic equivalence, the main judgment is whether the formal expression generated by autoformalization accurately reflects the corresponding natural language expression. Relevant indicators commonly used similarity measures including text similarity (e.g., BLEU, ROUGE, etc.), Poiroux et al. (2025) presents the symbolic computation index: BEqL and BEq +. As pointed out by Liu et al. (2025), BEqL can indicate the accuracy of various syntactic changes. This method can achieve the accuracy of various syntactic changes. Although it has a high false negative rate, the paper believes that its simplicity makes this indicator worth using. BEq+ introduces more complex policy combinations (such as simp, ring, etc.) on the basis of BEqL, further enhancing the ability to verify the accuracy of deep mathematical structures (such as algebraic identities). Lu et al. (2024a) mentioned that for LLM generated formal results, even if the BLEU index is high also don’t tell the formal result with the corresponding natural language expression is semantic equivalence, Alignment scoring is also proposed to better evaluate semantic equivalence.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Beyond mathematics, however, the core principles of autoformalization—translating informal or semi-formal language into precise, verifiable representations—hold broader significance. As LLMs are increasingly applied to tasks like algorithm design, code generation, and system specification Shi et al. (2023); Di et al. (2024); Zheng et al. (2025), ensuring that their outputs are not only syntactically correct but also semantically sound becomes critical. Autoformalization offers a promising mechanism for bridging this gap: by converting natural language specifications, pseudo-code, or even informal correctness claims into formal logic or verifiable code contracts, it enables rigorous checking of algorithmic behavior. In domains such as software engineering, security auditing, and scientific computing, this capability could substantially enhance the trustworthiness and reliability of LLM-generated artifacts. Viewed from this perspective, autoformalization is not merely a tool for mathematical proof—it is a foundational enabler for building LLM systems that can reason, code, and verify in an integrated and principled way.
@@
This survey provides a comprehensive and structured review of the recent advances in automatic formalization, focusing on the current applications in mathematical theorem proving and the possibility of combining automatic formalization with formal verification of large models to generate fully trustworthy outputs in the future. §2 shows a review of the historical and technological foundations of autoformalization. And we summarize notable open-source models and datasets that support current research in §3; In §4, we dissect the autoformalization workflow, covering key components such as data preprocessing, modeling strategies, postprocessing techniques, and evaluation protocols, and also dissect state of the art model approaches to mathematical theorem proving. In §5, we conduct a deep analysis of the involved mathematical problems from the mathematical problem domain, the mathematical problem difficulty level and the mathematical problem abstraction level. In §6, we explore the connection between autoformalization and the formal verification of large model output from multiple fields. In §7, we discuss major open challenges and promising future research directions. Finally, §8 compares our survey with existing related surveys, clarifying its unique perspective and contributions.
Table 2: Datasets/Benchmarks with code repos link and description
Dataset/Benchmark & Link Size Description
MiniF2F 488 Statements from olympiads as well as high-school and undergraduate maths classes.
LeanEuclid 173 Euclidean geometry problems manually formalized in Lean.
Math 12,500 12,500 challenging competition mathematics problems.
FORML4 17,137 Informalize theorems extracted from Mathlib 4.
ProofNet 374 The problems are drawn from popular undergraduate pure mathematics textbooks.
PutnamBench 657 Competition mathematics problems sourced from the William Lowell Putnam Mathematical Competition.
LeanDojo 122,517 Theorems,proofs,tactics,premises from mathlib4
GSM8K 8,792 High quality linguistically diverse grade school math word problems.
miniCTX 762 Theorems sourced from real Lean projects and textbooks.
arxiv2Formal 50 Theorems from arXiv papers formalized in Lean.
DeepSeek-Prover-V1 27,503 Lean4 proof data derived from high-school and undergraduate-level mathematical competition problems.
Lean-Workbook 13,517 Collection of mathematical problems and their proof information.
Lean-GitHub 20,446 Theorems with formal proofs and tactics from 2,133 files.
Compfiles 205 A collection of olympiad-style math problems and their solutions, formalized in Lean 4.
Herald 624,436 The dataset generated from Herald pipeline on Mathlib4, containing 580k valid statements and 44k NL-FL proofs.
ProverBench 325 A benchmark dataset containing 325 formalized problems to advance neural theorem proving research, including 15 from the prestigious AIME competitions (years24-25).
CombiBench 100 A manually produced benchmark, including 100 combinatorial mathematics problems of varying difficulty and knowledge levels.
STP_Lean_0320 3262558 A dataset including mathlib4 examples, LeanWorkbook proofs, and self-play conjecture proofs.
Leanabell-Prover-Formal-Statement 1,133,283 A dataset including released data by existing models and synthetic data from informal math problems.
IneqComp 75 A benchmark built from elementary inequalities through systematic transformations, including variable duplication, algebraic rewriting, and multi-step composition.
APE-Bench I 10928 The first large-scale, file-level benchmark for proof engineering, built from over 10,000 real-world commits in Mathlib4.
Formal Math 5560 A benchmark of 5,560 formally verified mathematical statements spanning diverse subdomains. This dataset is 22.8× larger than the widely used MiniF2F benchmark.
DeepTheorem 121000 It comprises 121K IMO-level informal theorems and proofs spanning diverse mathematical domains.
Numinamath 860000 A dataset of math competition problem-solution pairs, with each solution templated with Chain of Thought (CoT) reasoning.
Table 3: Availability of Informal/Formal Statements and Solutions Across Datasets
Dataset & Link Informal Statement Formal Statement Informal Solution Formal Solution
[…]
@@ 2.2 Autoformalization in LLM Output Formal Verification
To address this challenge, recent research has explored the use of autoformalization as a means of translating natural language outputs from LLMs into precise, verifiable representations using formal systems such as first-order logic Ryu et al. (2024) or arithmetic system such as Peano arithmeticPeano (1889). Balcan et al. (2025) is also a recent research which focus on the chain of thought by designing a formal definition verifier and training the model to learn the correctness of the reasoning sequence, aiming to ensure that each reasoning step of the chain of thought is valid. In this context, autoformalization acts as a bridge between the inherent ambiguity of natural language and the strict requirements of formal verification. By enabling symbolic reasoning over LLM outputs, autoformalization enhances correctness, improves reliability, and promotes the safe and trustworthy deployment of LLMs in tasks requiring rigorous reasoning and high assurance.
@@
3 Foundations of Progress: Techniques and Corpora
  • Formal Proof Gen: Using the reasoning ability of llm, the given natural language mathematical problem statement is automatically generated into formal proofs that can be verified by formal systems, and the corresponding grammar rules and solution logic are satisfied.
  • Data Construction: Through large-scale synthesis and generation, enhancing or filtering data, the bottleneck of data scarcity is broken and the performance of the model is improved. It aims to systematically extract, annotate, and structure the training data required for formal model learning from existing theorem libraries, proof scripts, or natural language mathematics. The goal is to provide high-quality, large-scale parallel prediction for downstream tasks, supporting supervised learning and reinforcement learning training.
  • Evaluation: Compilation of formal proof assistans does not mean that autoformalization is completely correct. In order to better verify the correctness of autoformalization, new methods or performance metrics are proposed.
  • Autoformalization: It involves the automatic conversion of mathematical statements expressed in natural language into formal language statements. This task requires the model to complete a representation that can be verified by a formal system under multiple constraints such as semantic consistency, contextual completeness, and symbolic accuracy, forming a key bridge between natural language and formal systems.
  • RAG: Given the current state of the proof, enhance the ability to retrieve potential premises and establish reflection between them. By retrieving semantically related previous proofs or structures from the theorem library, RAG assists the generator to construct effective reasoning trajectories more efficiently, and improves the accuracy and verifiability of generation.
Table 2 provides a comprehensive overview of the datasets used for theorem proving and automatic formalization, including the addresses, sizes, and brief descriptions of the datasets. Table 3 gives the support that the individual datasets contain for the general dataset elements. Includes informal problem statement, formal problem statement, informal solution, formal solution. The main task is to summarize the task categories supported by the datasets. Each element in Table 3 is described as follows:
  • Informal Statement:It refers to the description of mathematical problems expressed in natural language, usually close to the way of expression in mathematics textbooks or Olympiad questions, suitable for human understanding but not strictly following the grammar of formal language.
  • Formal Statement:Problem definitions accurately represented by formal languages (such as the syntax of proof assistants like Lean, Isabelle, and Coq) possess logical precision and verifiability, and can serve as objective theorems in formal systems.
  • Informal Solution:The problem-solving process or answer expressed in natural language or standard mathematical writing forms usually includes reasoning steps, auxiliary lemmas, etc., which are convenient for manual review but do not have verifiability in a formal system.
  • Formal Solution:The corresponding formal language code, which is usually checked by a theorem proving assistant, is the core component to achieve automatic verification and program aided proof.
@@ 4.1.2 Increase data quantity
Due to a large number of written by human still didn’t get the full use of formal language corpus, in order to solve this problem Wu et al. (2024b) collects almost all of the Lean4 repositories on GitHub, and makes suitable training data; From Mathlib4’s 10,000 historical real-world commits, Xin et al. (2025a) builds the first large-scale, file-level dataset. Both of these two methods increase the amount of data, and experiments show that training on such a large scale dataset, the model improves performance in a variety of mathematical problem domains.
In order to further expand the data, Liu et al. (2025c) extracts mathematical concepts from mathlib, and randomly selects two concepts to let the LLM generates informal problem statements, and uses the model with more parameters as the teacher model to generate the corresponding formal proofs, thus creating a new massively parallel corpus. In order to further expand the data, Liu et al. (2025c) also disproves the assumptions in the formal proof, so that the number of parallel corpora created in the end is nearly doubled. Using the final created parallel corpus, the student model is guided through reinforcement learning methods to iterate over its autoformalization capabilities.
@@ 4.2.2 LLM-based Autoformalization
In natural language machine translation related work, the quality of translation largely depends on the quality of parallel data between the two languages. But such parallel data are rare and difficult to manage, and back-translation is one of the most effective ways to improve the quality of translation. Azerbayev et al. (2023) uses OpenAI distillation of the Codex model of back translation, in order to improve the model of automatic formal ability. Jiang et al. (2024) They used GPT-4 to generate informal language data from an existing library of formal languages. They have empirically and analytically concluded that informal is much simpler than formal, so back-translation is part of the Autoformalization effort.Chan et al. (2025) explores three approaches to back-translation and constructs a high-quality, small-scale dataset, which outperforms MMA,ProofNet using only 1/150 of the number of tokens. First, it trains a single model to handle both directions of the translation task, it does not employ a teacher model and a student model as in the traditional practice: (1) Transform a batch of examples to informal language; (2) Using these synthetic informal language examples and formal language example pairs as labeled training data, the synthetic informal language examples are back-translated to formal language; (3)Compute the loss between the generated formal language examples and the original formal language examples; (4)Update the model weights via backpropagation. Experiments significantly reduce the evaluation loss after only a few hundred training steps, resulting in a significant improvement over the baseline; Another method uses a pretrained model with more parameters and stronger performance as a teacher model, provides an formal language example dataset, and uses the model to generate the corresponding synthetic informal language text, and then trains the student model to generate formal language according to the informal language, so that the ability of the small-scale student model approaches that of the large-scale teacher model.The last method captures the specific keyword parsing formal proof strategy through regular expressions, which is more interpretable than the above two methods, but due to the limitations of the format, only partially simple translations can be generated, and it is difficult to surpass the above complex back-translation methods.
@@ 4.2.4 Evaluatation
In the Autoformalization workflow, syntactic accuracy and semantic equivalence at the evaluation level are key factors to ensure the success of the autoformalization process. Syntactic accuracy refers to the ability to correctly translate into formal language, that is, whether the final formal result can pass the type check of Lean, Coq and other environments. It is also possible to extend first-order logic formal languages with syntax close to natural languages but more mathematically expressive. Zhou and Zhang (2025) uses a LLM to translate the natural language problem solution steps into a kind of designed formal language SimpleMath, and take a data-oriented approach to structure the formalized solution into a solution Graph, which shows the dependencies of each reasoning step. When verifying the correctness of a reasoning step, it does not need to input all constraints into the verification model, but selects the most relevant premise to the current reasoning conclusion, which greatly reduces the input length and improves the efficiency. Compared with formal languages such as Lean and Coq, it is more flexible. The type checking of Lean, Coq and other environments does not mean that it is completely correct, because the formalized results also need to be semantically equivalent to the corresponding natural language version, so it is equally important to evaluate the semantic equivalence. The purpose of grammatical accuracy is to ensure the result of autoformalization through the formal language proof assistant’s type checking mechanism, which mainly involves two aspects: 1) at the level of code generation: the generated formal language code has no grammatical errors; 2) At the level of mathematical reasoning: the generated formal language code has a correct logical chain. The most common way to assess accuracy is to compare automatically generated formal expressions with the results of expert manual formalization. Such evaluation often rely on recall metrics or manual inspections(Wu et al. (2022) Agrawal et al. (2022), Azerbayev et al. (2023), Jiang et al. (2023a))to judge the consistency of the two in mathematical content. Although it is right, this method is time-consuming and subjective, so it is commonly used to calibrate automatic evaluate tools. For the accuracy of mathematical reasoning, such as whether the model corresponds to the specific meaning of each variable, or whether the proof strategy used in the proof process is correct, For the former, Li et al. (2024a)suggests the method of symbolic equivalence . Split statements into premises and conclusions (assuming that premises do not contradict themselves) When the premises and conclusions of the formal results to be verified are respectively equivalent to the corresponding correct premises and conclusions by ATP proof, it is explained that the model corresponds to the specific meaning of each variable. For the latter Zheng et al. (2024) leverage the feedback to evaluate whether the proof strategy is correct and the tool correction mechanism is used to resolve it. For semantic equivalence, the main judgment is whether the formal expression generated by autoformalization accurately reflects the corresponding natural language expression. Relevant indicators commonly used similarity measures including text similarity (e.g., BLEU, ROUGE, etc.), Poiroux et al. (2025) presents the symbolic computation index: BEqL and BEq +. As pointed out by Liu et al. (2025b), BEqL can indicate the accuracy of various syntactic changes. This method can achieve the accuracy of various syntactic changes. Although it has a high false negative rate, the paper believes that its simplicity makes this indicator worth using. BEq+ introduces more complex policy combinations (such as simp, ring, etc.) on the basis of BEqL, further enhancing the ability to verify the accuracy of deep mathematical structures (such as algebraic identities). Lu et al. (2024a) mentioned that for LLM generated formal results, even if the BLEU index is high also don’t tell the formal result with the corresponding natural language expression is semantic equivalence, Alignment scoring is also proposed to better evaluate semantic equivalence. Recent studies have also proposed non-numerical metrics to verify semantic equivalence: Xuejun et al. (2025) proposes the LeanScorer method, which uses a LLM to decompose natural language statements into premises and conclusions. Then, another LLM is used to score the premises and conclusions corresponding to the formal statements generated by the model respectively. Finally, the scoring results are integrated into numerical scores to test the semantic equivalence of the formal statements;Raza and Milic-Frayling (2025) proposes a new semantic verification method, semantic self-verification, which can solve the key challenge of combining language models with the rigor of logic solvers without training a model: how to more accurately translate reasoning problems in natural language into a formal language for solvers. Raza and Milic-Frayling (2025) generates the solution program through LLM, and then generates specific positive and negative examples to verify whether the formal results meet the two kinds of test instances to judge the semantic correctness of the formal results. It also uses examples to constrain, so that LLM can better modify the failed formal results.